# R-SCoRe: Revisiting Scene Coordinate Regression for Robust Large-Scale Visual Localization

### v1

@@ 4.3 Implementation Details
Network Architecture. We adopt the MLP architecture and position decoder from GLACE [74] with expansion ratio $m=2$ for the MLP, and $k=50$ clusters for the position decoder. The network width of the MLP is set to 768 for simpler scenes and 1280 for more complex scenes.
@@
Testing. At test time, we retrieve the 10 nearest training images with NetVLAD [1]. The global encoding and retrieval features for training images are precomputed and compressed using Product Quantization [30]. For final pose estimation, we utilize PoseLib [35] with a maximum reprojection error of 10 pixels and up to 10,000 RANSAC iterations. The total inference time for each query image is approximately 140 to 270 ms with Intel i7-9700K CPU and NVIDIA RTX 2080 GPU.
@@ 4.4 Ablation Study Results
In our ablation studies, we investigate the impact of the different components in R-SCoRe. Since the Hyundai Department Store dataset provides a validation split, we conduct most ablation studies on this split. In contrast, the Aachen Day-Night dataset does not provide a validation split and, thus, we evaluate on the test set.
@@
Global Encoding. Retrieving global encodings from training images avoids the domain gap in global encoding. Better retrieval method and multiple hypotheses verification can help resolve ambiguities. As shown in Tab. 5, without retraining, utilizing multiple global hypotheses at test time (+ Multi Hypotheses) results in a direct performance improvement in complex scenes. The performance improves significantly, once we incorporate our covisibility graph-based data augmentation during training. In particular, we replace isotropic Gaussian noise [74] with our more precise covisibility-based technique (+ Covis Augmentation). Finally, learning the global encoding directly from the covisibility graph (+ Covis Encoding) reduces the interference between non-covisible training images and thereby facilitates implicit triangulation, especially in indoor scenes with significant ambiguity.
@@ 5 Conclusion
In this work, we revisited scene coordinate regression (SCR) methods for robust visual localization in large-scale, complex environments. We analyzed the design principles of input encoding and training strategies, identifying several areas for enhancement. Our proposed R-SCoRe includes a covisibility graph-based global encoding learning and data augmentation strategy, a depth-adjusted reprojection loss to improve the implicit triangulation, and also other improvements including better architecture and local feature. Our contributions advance the state-of-the-art in SCR and demonstrate that SCR-based localization methods can achieve competitive performance in large-scale applications. While operating at comparably very small map sizes, R-SCoRe trails the state-of-the-art FM-based localization methods only at the strictest error thresholds. Given the relatively small history of SCR, we are positive the accuracy gap can be closed completely in the near future.

### v2

@@ 4.3 Implementation Details
Most hyperparameters follow default values [75] and extensive tuning is not performed, as we empirically find the approach remains robust within a reasonable range, apart from the trade-off between network size and performance.
@@
Network Architecture. We adopt the MLP architecture and position decoder from GLACE [75] with expansion ratio $m=2$ for the MLP, and 50 clusters for the position decoder. With MLP width $w\!=\!256\left\lceil\sqrt{n/1000}\right\rceil$ for $n$ training images, we scale the parameter count proportionally.
@@
Testing. At test time, we retrieve the 10 nearest training images with NetVLAD [1]. The global encoding and retrieval features for training images are precomputed and compressed using Product Quantization [30]. For final pose estimation, we utilize PoseLib [35] with a maximum reprojection error of 10 pixels and up to 10,000 RANSAC iterations. On our PC (NVIDIA RTX 2080 GPU & Intel i7-9700K CPU), the average inference time for a $640\!\times\!480$ query image is 140 to 270 ms in total.
  • Global: NetVLAD (20ms), Retrieval ($<$1ms)
  • Local: LoFTR (7ms) or DeDoDe (50ms)
  • MLP: w = 768 (70ms) or 1280 (160ms)
  • Pose Solving: 40ms
@@ 4.4 Ablation Study Results
In our ablation studies, we investigate the impact of the different components in R-SCoRe. We evaluate on the validation split for the Hyundai Department Store dataset. Since the Aachen Day-Night dataset does not provide a validation split, we evaluate on the test set.
@@
Global Encoding. Retrieving global encodings from training images avoids the domain gap. Better retrieval method and multiple hypotheses verification can help resolve ambiguities. Without retraining (Tab. 5), utilizing multiple global hypotheses at test time (+ Multi Hypotheses) results in a direct performance improvement in complex scenes. The performance improves significantly, once we incorporate our covisibility graph-based data augmentation during training. In particular, we replace isotropic Gaussian noise [75] with our more precise covisibility-based technique (+ Covis Augmentation). Finally, learning the global encoding directly from the covisibility graph (+ Covis Encoding) reduces the interference between non-covisible training images and thereby facilitates implicit triangulation, especially in indoor scenes with significant ambiguity.
@@ 5 Conclusion
In this work, we revisited scene coordinate regression (SCR) methods for robust visual localization in large-scale, complex environments. We analyzed the design principles of input encoding and training strategies, identifying several areas for enhancement. Our proposed R-SCoRe includes a covisibility graph-based global encoding learning and data augmentation strategy, a depth-adjusted reprojection loss to improve the implicit triangulation, and also other improvements including better architecture and local feature. Our contributions advance the state-of-the-art in SCR and demonstrate that SCR-based localization methods can achieve competitive performance in large-scale applications. While operating at comparably very small map sizes, R-SCoRe trails the state-of-the-art FM-based localization methods only at the strictest error thresholds. Although out-of-distribution generalization remains challenging, and gaps persist in handling extreme cases, given the relatively small history of SCR, we are positive the accuracy gap can be closed completely in the near future.